# Classification of abelian Schur groups I

### v1

@@ 7.2. Nondense $S$-rings
If $A\times P_{1}=G$, then we are done by Eq. (8). So we may assume that $A\times P_{1}<G$. If $P\leq\operatorname{rad}(Z)$ for every $Z\in\mathcal{S}(\mathcal{A})_{G\setminus(A\times P_{1})}$, then $\mathcal{A}$ is the nontrivial $U/L=(A\times P_{1})/P_{1}$-wreath product and $U/L$ satisfies Condition $(1)$.
@@
Due to the above discussion, we may assume that there is $Z\in\mathcal{S}(\mathcal{A})_{G\setminus(A\times P_{1})}$ with $P\nleq\operatorname{rad}(Z)$. Then Statement $(2)$ of Lemma 7.2 holds for $Z$, i.e $\langle(Z^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$. Eq. (8) and $P_{1}\nleq D$ imply that $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $H$ and hence $\langle(Z^{[p]})_{H}\rangle=A$. The latter equality yields that $Z^{[p]}\subseteq A\times D$. The set $Z^{[p]}$ is an $\mathcal{A}$-set by Lemma 2.4 and consequently $\langle Z^{[p]}\rangle$ is an $\mathcal{A}$-subgroup of $A\times D$. However, $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $A\times D$ by Eq. (8) and $P_{1}\nleq D$. Therefore $\langle Z^{[p]}\rangle=A$. From the latter equality it follows that $Z\cap(A\times P_{1})\neq\varnothing$. As $A\times P_{1}$ is an $\mathcal{A}$-subgroup, we obtain $Z\subseteq A\times P_{1}$, a contradiction to the definition of $Z$.
@@
Case 2. Suppose that Statement $(1)$ or Statement $(2)$ of Lemma 7.2 holds for every basic set of $\mathcal{A}$ outside $D$. Let $N$ be the $\mathcal{A}$-subgroup generated by all basic sets $X$ of $\mathcal{A}$ such that $P\nleq\operatorname{rad}(X)$. Clearly, $N\geq P_{1}$. We are done if $|N|$ is not divisible by $4$. Indeed, then $\mathcal{A}$ is the nontrivial $U/L=N/P_{1}$-wreath product and $U/L$ satisfies Condition $(3)$ from Proposition 7.3. Assume that $|N|$ is divisible by $4$ or, equivalently, $N\geq H$. Then there is $Y\in\mathcal{S}(\mathcal{A})_{G\setminus D}$ such that $P\nleq\operatorname{rad}(Y)$. Lemma 7.2 implies that $A=\langle(Y^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$. As $H$ is not an $\mathcal{A}$-subgroup, $|A|=2$.
@@
Since $N\geq H$, there is $Z\in\mathcal{S}(\mathcal{A})_{N}$ such that $P\nleq\operatorname{rad}(Z)$ and $Z_{H}\nsubseteq A$. Note that $Z\nsubseteq H$ and hence
@@
Let $V=\langle Z\rangle$. The $S$-ring $\mathcal{A}_{V/A}$ is an $S$-ring over a cyclic group $V/A\cong C_{2p^{i}}$, $i\in\{1,\ldots,k\}$. Due to Eq. (9) and the assumption that $H$ is not an $\mathcal{A}$-subgroup, we have $|\operatorname{rad}(Z^{\pi})|=1$, where $\pi$ is the canonical epimorphism from $G$ to $G/A$. So $|\operatorname{rad}(\mathcal{A}_{V/A})|=1$. Note that $\mathcal{A}_{V/A}$ can not be cyclotomic or tensor product of two $S$-rings because $Z$ contains elements of even as well as odd orders. Therefore $\mathcal{A}_{V/A}$ is trivial by Lemma 6.2.
@@
for all $z\in Z_{D}\setminus\{e\}\neq\varnothing$ and $h\in H$ such that $hz\in Z$.
Observe that $Z\cap(A\times P^{\#})\neq\varnothing$ because $\mathcal{A}_{V/A}=\mathcal{T}_{V/A}$. Eq. (11) yields that $P^{\#}\subseteq Z$ or $aP^{\#}\subseteq Z$. If the both of these two inclusions hold, then $A\leq\operatorname{rad}(Z)$ by Lemma 2.1. Further, we may assume that exactly one of the above inclusions hold. Then $P^{\#}\subseteq Z$ or $P^{\#}\subseteq aZ$. Let $Z^{\prime}\in\{Z,aZ\}$ such that $P^{\#}\subseteq Z^{\prime}$. From Eqs. (10) and (11) it follows that $Z^{\prime}\setminus P\neq\varnothing$ and $P\leq\operatorname{rad}(Z^{\prime}\setminus P)$. Therefore $|\operatorname{rad}(Z)|=1$ and $Z^{\prime}=\langle Z\rangle^{\#}$ by Lemma 2.2. We conclude that Statement $(3)$ of Lemma 7.2 holds for $Z^{\prime}$, a contradiction to the assumption of this case. Thus, $A\leq\operatorname{rad}(Z)$. Due to triviality of $\mathcal{A}_{V/A}$, we obtain
@@
If $V=G$, then due to Eq. (12), $\mathcal{A}$ is the $U/L=A/A$-wreath product and $U/L$ satisfies Condition $(1)$. Further, we assume that $V<G$. If there exists $T\in\mathcal{S}(\mathcal{A})_{G\setminus V}$ such that $T\subseteq D$, then $\langle T\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $D$. So $\langle T\rangle\cap V\neq\varnothing$ and hence $V\subseteq\langle T\rangle\leq D$, a contradiction to the definition of $Z$. Therefore every basic set outside $V$ is not a subset of $D$ and hence Statement $(1)$ or $(2)$ of Lemma 7.2 holds for every such basic set. If Statement $(1)$ of Lemma 7.2 holds for every basic set $X$ outside $V$, i.e. $P\leq\operatorname{rad}(X)$, then $V=P_{1}\leq\operatorname{rad}(X)$. This implies that $\mathcal{A}$ is the $U/L=V/V$-wreath product and $U/L$ satisfies Condition $(1)$. If Statement $(2)$ of Lemma 7.2 holds for some $T\in\mathcal{S}(\mathcal{A})_{G\setminus V}$, then $\langle(T^{[p]})_{H}\rangle=A$ because $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $H$. Together with Lemma 2.4, this yields that $W=\langle T^{[p]}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $A\times D$. Since $T$ lies outside $V$, we have $(T^{[p]})_{D}\neq\{e\}$ and hence $|W|$ is divisible by $p$. Therefore $W\cap Z\neq\varnothing$. Thus,
@@
The case when $D$ is not an $\mathcal{A}$-subgroup is dual to the case when $H$ is not an $\mathcal{A}$-subgroup in the sense of the duality theory of $S$-rings over an abelian group [31] (see also [7, Section 2.2]). The proposition below immediately follows from Lemma 4.1 and Proposition 7.3.

### v3

@@ 7.2. Nondense $S$-rings
by Lemma 3.2(2). If $A\times P_{1}=G$, then $\mathcal{A}$ is a nontrivial tensor product by Eq. (8) and we are done. So we may assume further that $A\times P_{1}<G$.
If $P\leq\operatorname{rad}(Y)$ for every $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(A\times P_{1})}$, then $\mathcal{A}$ is the nontrivial $U/L=(A\times P_{1})/P_{1}$-wreath product and $U/L$ satisfies Condition $(1)$ from Theorem 7.3 as desired.
@@
If there is $Y\in\mathcal{S}(\mathcal{A})_{G\setminus(A\times P_{1})}$ such that $P\nleq\operatorname{rad}(Y)$, then Statement $(2)$ of Lemma 7.2 holds for $Y$, i.e $\langle(Y^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$. Eq. (8) and $P_{1}\nleq D$ imply that $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $H$ and hence $\langle(Y^{[p]})_{H}\rangle=A$. The latter equality yields that $Y^{[p]}\subseteq A\times D$. The set $Y^{[p]}$ is an $\mathcal{A}$-set by Lemma 2.4 and consequently $\langle Y^{[p]}\rangle$ is an $\mathcal{A}$-subgroup of $A\times D$. However, $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $A\times D$ by Eq. (8) and $P_{1}\nleq D$. Therefore $\langle Y^{[p]}\rangle=A$. From the latter equality, it follows that $Y\cap(A\times P_{1})\neq\varnothing$. As $A\times P_{1}$ is an $\mathcal{A}$-subgroup, we obtain $Y\subseteq A\times P_{1}$ in contrast to the definition of $Y$.
@@
Case 2. Suppose that Statement $(1)$ or Statement $(2)$ of Lemma 7.2 holds for every basic set of $\mathcal{A}$ outside $D$. Clearly, $P\nleq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})$ such that $X\cap P\neq\varnothing$. So $N\geq P_{1}$, where $N$ is the $\mathcal{A}$-subgroup generated by all basic sets $X$ of $\mathcal{A}$ such that $P\nleq\operatorname{rad}(X)$.
We may assume that $|N|$ is divisible by $4$: otherwise $\mathcal{A}$ is the nontrivial $U/L=N/P_{1}$-wreath product and $4\nmid|U|$ and $|L|\neq 2$ as required. Assume that $|N|$ is divisible by $4$ or, equivalently, $N\geq H$. Then there is $Y\in\mathcal{S}(\mathcal{A})$ outside $D$ such that $P\nleq\operatorname{rad}(Y)$ because otherwise $N\leq D$ and hence $N\ngeq H$. Lemma 7.2 implies that $A=\langle(Y^{[p]})_{H}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $H$. As $H$ is not an $\mathcal{A}$-subgroup, $|A|=2$.
@@
Since $N\geq H$, there is $Z\in\mathcal{S}(\mathcal{A})_{N}$ such that $P\nleq\operatorname{rad}(Z)$ and $Z_{H}\nsubseteq A$ because otherwise $N\leq A\times D$ and hence $N\ngeq H$. Note that $Z\nsubseteq H$ and hence
@@
Let $V=\langle Z\rangle$. The $S$-ring $\mathcal{A}_{V/A}$ is an $S$-ring over a cyclic group $V/A\cong C_{2p^{i}}$, $i\in\{1,\ldots,k\}$. Every basic set of a cyclotomic $S$-ring or a nontrivial tensor product over $V/A$ is regular. However, $Z$ and hence $Z^{\pi}$, where $\pi$ is the canonical epimorphism from $G$ to $G/A$, are nonregular because $Z_{D}\setminus\{e\}\neq\varnothing$ and $Z_{H}\nsubseteq A$. Therefore $\mathcal{A}_{V/A}$ can not be cyclotomic or a nontrivial tensor product. Lemma 6.2 implies that $\mathcal{A}$ is trivial or $|\operatorname{rad}(\mathcal{A}_{V/A})|>1$. Observe that $|\operatorname{rad}(Z^{\pi})|=1$ by Eq. (9) and the assumption that $H$ is not an $\mathcal{A}$-subgroup and hence the latter is impossible. Thus, $\mathcal{A}_{V/A}$ is trivial.
@@
for all $z\in Z_{D}\setminus\{e\}$ and $h\in H$ such that $hz\in Z$.
Observe that $Z\cap(A\times P^{\#})\neq\varnothing$ because $\mathcal{A}_{V/A}=\mathcal{T}_{V/A}$. Eq. (11) yields that $P^{\#}\subseteq Z$ or $aP^{\#}\subseteq Z$. If both of these two inclusions hold, then $A\leq\operatorname{rad}(Z)$ by Lemma 2.1. Thus, we may assume that exactly one of the above inclusions holds. Then $P^{\#}\subseteq Z$ or $P^{\#}\subseteq aZ$. Let $Z^{\prime}\in\{Z,aZ\}$ be such that $P^{\#}\subseteq Z^{\prime}$. From Eqs. (10) and (11), it follows that $Z^{\prime}\setminus P\neq\varnothing$ and $P\leq\operatorname{rad}(Z^{\prime}\setminus P)$. Therefore $|\operatorname{rad}(Z)|=1$ and $Z^{\prime}=\langle Z\rangle^{\#}$ by Lemma 2.2. We conclude that Statement $(3)$ of Lemma 7.2 holds for $Z^{\prime}$, a contradiction to the assumption of this case. Thus, $A\leq\operatorname{rad}(Z)$. Due to triviality of $\mathcal{A}_{V/A}$, we obtain
@@
If $V=G$, then due to Eq. (12), $\mathcal{A}$ is the $U/L=A/A$-wreath product and $U/L$ satisfies Condition $(1)$ from Theorem 7.3. Further, we assume that $V<G$. If there exists $T\in\mathcal{S}(\mathcal{A})_{G\setminus V}$ such that $T\subseteq D$, then $\langle T\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $D$. So $\langle T\rangle\cap V\neq\varnothing$ and hence $V\subseteq\langle T\rangle\leq D$, a contradiction to the definition of $Z$. Therefore every basic set outside $V$ is not a subset of $D$ and hence Statement $(1)$ or $(2)$ of Lemma 7.2 holds for every such basic set. If Statement $(1)$ of Lemma 7.2 holds for every basic set $X$ outside $V$, i.e. $P\leq\operatorname{rad}(X)$, then $V=P_{1}\leq\operatorname{rad}(X)$. This implies that $\mathcal{A}$ is the $U/L=V/V$-wreath product and $U/L$ satisfies Condition $(1)$ from Theorem 7.3. If Statement $(2)$ of Lemma 7.2 holds for some $T\in\mathcal{S}(\mathcal{A})_{G\setminus V}$, then $\langle(T^{[p]})_{H}\rangle=A$ because $A$ is a unique nontrivial $\mathcal{A}$-subgroup of $H$. Together with Lemma 2.4, this yields that $W=\langle T^{[p]}\rangle$ is a nontrivial $\mathcal{A}$-subgroup of $A\times D$. Since $T$ lies outside $V$, we have $(T^{[p]})_{D}\neq\{e\}$ and hence $|W|$ is divisible by $p$. Therefore $W\cap Z\neq\varnothing$. Thus,
@@
If $H$ is not an $\mathcal{A}$-subgroup and $\mathcal{A}$ is a nontrivial $U/L$-wreath product for an $\mathcal{A}$-section $U/L$ such that $4\nmid|U|$ and $|L|\neq 2$, then Theorem 7.3 holds.
Proof.
Note that $U$ is cyclic because $4\nmid|U|$. If $|\operatorname{rad}(\mathcal{A}_{U})|=1$, then $U/L$ satisfies Condition $(3)$ from Theorem 7.3 and we are done. From now on throughout the proof of this proposition, we assume that
(13) $|\operatorname{rad}(\mathcal{A}_{U})|>1.$
Suppose that $|U|$ is odd. Then $|U|=pq$ or $|U|$ is a $p$-power. In the former case, Eq. (13) and Lemma 6.12 imply that $\mathcal{A}_{U}=\mathcal{A}_{L}\wr\mathcal{A}_{U/L}$. So $\mathcal{A}=\mathcal{A}_{L}\wr\mathcal{A}_{G/L}$ and Theorem 7.3 holds as required. In the latter one, $\mathcal{A}_{U}$ is the nontrivial $U_{1}/L_{1}$-wreath product for an $\mathcal{A}_{U}$-section $U_{1}/L_{1}$ of $U$ such that $L_{1}$ is the least nontrivial $\mathcal{A}_{U}$-subgroup and $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$ by Eq. (13) and Lemma 6.10. By the definitions of $U_{1}$ and $L_{1}$, we have $U_{1}\leq U$ and $L_{1}\leq L$. This implies that $\mathcal{A}$ is the $U_{1}/L_{1}$-wreath product and $U_{1}/L_{1}$ satisfies Condition $(3)$ from Theorem 7.3 as required.
Now suppose that $|U|$ is even. Then $|U|=2m$, where $m$ is odd, because $4\nmid|U|$. Let $A$ and $W$ be the Sylow $2$-subgroup and the Hall $2^{\prime}$-subgroup of $U$, respectively. If $|W|=1$, then $|L|=|U|=2$, a contradiction to the assumption $|L|\neq 2$ of the lemma. In what follows, we assume that
$|W|>1.$
The least $\mathcal{A}_{U}$-subgroup containing $A$ and a maximal $\mathcal{A}_{U}$-subgroup of $W$ are denoted by $A_{1}$ and $W_{1}$, respectively. Assume that $|W_{1}|=1$. Then $\mathcal{A}_{U}$ is trivial or $\mathcal{A}_{U}=\mathcal{A}_{A_{1}}\wr\mathcal{A}_{U/A_{1}}$ by Lemma 5.3 applied to $\mathcal{A}=\mathcal{A}_{U}$, $H_{1}=W_{1}$, and $P_{1}=A_{1}$. In the former case, we obtain a contradiction to Eq. (13). In the latter one, $A_{1}$ is the least nontrivial $\mathcal{A}_{U}$-subgroup because otherwise $A_{1}$ has a nontrivial $\mathcal{A}_{U}$-subgroup of odd order and hence $|W_{1}|>1$, a contradiction to the assumption $|W_{1}|=1$. So $A_{1}\leq L$. Together with $\mathcal{A}=\mathcal{A}_{U}\wr_{U/L}\mathcal{A}_{G/L}$ and $\mathcal{A}_{U}=\mathcal{A}_{A_{1}}\wr\mathcal{A}_{U/A_{1}}$, this yields that $\mathcal{A}=\mathcal{A}_{A_{1}}\wr\mathcal{A}_{G/A_{1}}$ and Theorem 7.3 holds as desired. In what follows, we assume that
$|W_{1}|>1.$
Lemma 7.6.
With the above notation, Theorem 7.3 holds unless $L\geq M$ for some minimal nontrivial $\mathcal{A}_{W_{1}}$-subgroup $M$.
Proof.
If $|L\cap W_{1}|>1$, then $L\cap W_{1}$ being an $\mathcal{A}_{W_{1}}$-subgroup contains a minimal nontrivial $\mathcal{A}_{W_{1}}$-subgroup and we are done. So we may assume that $|L\cap W_{1}|=1$. Since $|L|\neq 2$ (the assumption of the proposition) and $|W_{1}|>1$, this can happen only if $n=pq$, $U=W_{1}\times L$, $|W_{1}|$ is prime, $|L|$ is twice prime, and $L$ does not have a nontrivial $\mathcal{A}_{L}$-subgroup of odd order. Lemma 6.12 applied to $\mathcal{A}_{L}$ implies that
$\mathcal{A}_{L}=\mathcal{T}_{L}~\text{or}~\mathcal{A}_{L}=\mathcal{A}_{L_{0}}\wr\mathcal{A}_{L/L_{0}}$
for an $\mathcal{A}_{L}$-subgroup $L_{0}$ of order $2$.
In the former case, Lemma 6.2 applied to $\mathcal{A}_{U}$ yields that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\otimes\mathcal{T}_{L}$. Therefore $|\operatorname{rad}(\mathcal{A}_{U})|=1$, a contradiction to Eq. (13). In the latter one, Eq. (13) and Lemma 6.2 applied to $\mathcal{A}_{U}$ yield that $\mathcal{A}_{U}$ is the nontrivial $(W_{1}L_{0})/L_{0}$-wreath product. As $L_{0}\leq L$ and $\mathcal{A}$ is the $U/L$-wreath product, the $S$-ring $\mathcal{A}$ is the nontrivial $(W_{1}L_{0})/L_{0}$-wreath product. Since $|L_{0}|=2$, we have $\mathcal{A}_{L_{0}}=\mathbb{Z}L_{0}$. So $\mathcal{A}_{W_{1}L_{0}}=\mathcal{A}_{W_{1}}\otimes\mathbb{Z}L_{0}$ by Lemma 3.2(2). Therefore $\mathcal{A}_{L_{0}}$ is $\otimes$-complemented in $\mathcal{A}_{W_{1}L_{0}}$. Thus, the section $(W_{1}L_{0})/L_{0}$ satisfies Condition $(2)$ from Theorem 7.3 and we are done. ∎
By Lemma 7.6, we may assume till the end of the proof that
$L\geq M$
for some minimal nontrivial $\mathcal{A}_{W_{1}}$-subgroup $M$. One can see that $\mathcal{A}_{W_{1}}$ is the $V/M$-wreath product (possibly, trivial), where $V$ is an $\mathcal{A}_{W_{1}}$-subgroup with $|\operatorname{rad}(\mathcal{A}_{V})|=1$. Indeed, this follows from Lemma 6.2 and Lemma 6.12 if $|W_{1}|=pq$ and from Lemma 6.2 and Lemma 6.10 if $|W_{1}|$ is a $p$-power. Lemma 5.4 implies that
Let $W_{1}A_{1}=U$. If $W_{1}\cap A_{1}$ is trivial, then $\mathcal{A}_{A_{1}}$ is primitive and hence $\mathcal{A}_{A_{1}}=\mathcal{T}_{A_{1}}$ by Lemma 6.1. Eq. (14) implies that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\otimes\mathcal{A}_{A_{1}}$ and consequently $\mathcal{A}_{U}$ is the $U_{1}/L_{1}=(VA_{1})/M$-wreath product (possibly, trivial). As $M\leq L$, we conclude that $\mathcal{A}$ is the $U_{1}/L_{1}$-wreath product. Since $|\operatorname{rad}(\mathcal{A}_{V})|=1$ and
we obtain $|\operatorname{rad}(\mathcal{A}_{U_{1}})|=1$. Thus, $U_{1}/L_{1}$ satisfies Condition $(3)$ from Theorem 7.3 and we are done.
Now let $W_{1}A_{1}<U$. This implies that
$W_{1}<W.$
In this case, $|W_{1}|$ is a $p$-power and hence $M$ is the unique least $\mathcal{A}_{W_{1}}$-subgroup. Lemma 5.3 applied to $\mathcal{A}=\mathcal{A}_{U}$, $H_{1}=W_{1}$, and $P_{1}=A_{1}$ yields that $\mathcal{A}_{U}=\mathcal{A}_{W_{1}}\wr\mathcal{A}_{U/W_{1}}$ or
In the former case, $\mathcal{A}$ is the nontrivial $V/M$-wreath product because $\mathcal{A}_{W_{1}}=\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M}$ and $M\leq L$. Since $|\operatorname{rad}(\mathcal{A}_{V})|=1$, the section $V/M$ satisfies Condition $(3)$ from Theorem 7.3 and we are done.
Suppose that Eq. (15) holds. If $\mathcal{A}_{A_{1}}\neq\mathcal{T}_{A_{1}}$, then $A_{1}$ has a nontrivial $\mathcal{A}_{A_{1}}$-subgroup by Lemma 6.1 applied to $A_{1}$. This $\mathcal{A}_{A_{1}}$-subgroup is of odd order because $A_{1}$ is the least $\mathcal{A}_{U}$-subgroup containing $A$ and hence $|A_{1}\cap W_{1}|>1$. In view of Eqs. (14) and (15), we conclude that $\mathcal{A}_{U}$ is the nontrivial $W_{1}/(W_{1}\cap A_{1})$-wreath product. Since $M$ is the unique least $\mathcal{A}_{W_{1}}$-subgroup, we have $M\leq A_{1}\cap W_{1}$ and consequently $\mathcal{A}_{U}$ is the $W_{1}/M$-wreath product. As $\mathcal{A}_{W_{1}}=\mathcal{A}_{V}\wr_{V/M}\mathcal{A}_{W_{1}/M}$, we conclude that $\mathcal{A}_{U}$ is the nontrivial $V/M$-wreath product. Observe that $\mathcal{A}$ is the $V/M$-wreath product because $M\leq L$. The section $V/M$ satisfies Condition $(3)$ from Theorem 7.3 because $|\operatorname{rad}(\mathcal{A}_{V})|=1$ and we are done.
(16) $\mathcal{A}_{W_{1}A_{1}}=\mathcal{A}_{W_{1}}\otimes\mathcal{T}_{A_{1}}.$
If $A_{1}\neq A$, then $|W_{1}A_{1}|=2pq$, $U=W_{1}A_{1}$, and $L\in\{W_{1},A_{1}\}$. So $U/L$ satisfies Condition $(2)$ from Theorem 7.3 by Eq. (16) and we are done. Further, we assume that $A_{1}=A$.
Lemma 7.7.
With the above notation, let $X\in\mathcal{S}(\mathcal{A})$ be such that $X_{D}\nsubseteq W_{1}$. Then $A\leq\operatorname{rad}(X)$.
Proof.
If $X\subseteq U$, then $X\subseteq U\setminus(W_{1}A_{1})$. So $\operatorname{rad}(X)\geq A_{1}\geq A$ by Eq. (15). In the sequel, we assume that $X\nsubseteq U$. This implies that $L\leq\operatorname{rad}(X)$ because $\mathcal{A}$ is the $U/L$-wreath product. Let $R=\langle X\rangle$. Since $X_{D}\nsubseteq W_{1}$, we obtain
$R\cap(W\setminus W_{1})\neq\varnothing.$
Together with Eq. (15), this implies that $R\geq A_{1}\geq A$. The image of $Y\subseteq G$ under the canonical epimorphism from $G$ to $G/\operatorname{rad}(X)$ is denoted by $\mkern 1.5mu\overline{\mkern-1.5muY\mkern-1.5mu}\mkern 1.5mu$. Clearly, $|\operatorname{rad}(\mkern 1.5mu\overline{\mkern-1.5muX\mkern-1.5mu}\mkern 1.5mu)|=1$ and hence $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable.
Assume the contrary that $A\nleq\operatorname{rad}(X)$. Then $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu\geq\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Due to Eq. (15), we have $\operatorname{rad}(X)\cap D\leq W_{1}$ and
for every $Y\in\mathcal{S}(\mathcal{A})_{R}$ inside $R\cap(U\setminus(W_{1}A))$. Observe that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is nontrivial because $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$ is a proper nontrivial $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$-subgroup.
Suppose that $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is cyclic. Then $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is isomorphic to $C_{p^{i}}$ or $C_{2p^{i}}$, $i\geq 1$. As $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable, we conclude that $\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu})$ is trivial. So $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is cyclotomic or a nontrivial tensor product by Lemma 6.2. However, every basic set of such $S$-ring over $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ has trivial radical which contradicts to Eq. (17).
Now suppose that $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is noncyclic. Then $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu\cong E_{4}\times C_{p^{i}}$, $i\geq 1$. In view of Eq. (17), the Hall $2^{\prime}$-subgroup of $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu$ is not an $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$-subgroup. So the Sylow $2$-subgroup of $\widehat{G}$ is not an $\widehat{\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}}-$subgroup by Lemma 4.1(1). Since $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is indecomposable and nontrivial, so is $\widehat{\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}}$. Therefore $\widehat{\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}}$ is a nontrivial tensor product by Proposition 7.4 applied to $\widehat{\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}}$. Thus,
for some $\mathcal{A}_{R}$-subgroups $R_{1}$ and $R_{2}$ by Lemma 4.1(3). As $\operatorname{rad}(X)\geq L$ and $|L|\neq 2$, we may assume without loss of generality that $|\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}|$ is a power of $2$. Due to Eqs. (17) and (18), $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{1}\ngeq\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Therefore $\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}$ is a cyclic group containing $\mkern 1.5mu\overline{\mkern-1.5muA\mkern-1.5mu}\mkern 1.5mu$. Lemma 6.2 and Eq. (17) yield that $|\operatorname{rad}(\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu_{2}})|\neq 1$. Together with Eq. (18), this implies that $\mathcal{A}_{\mkern 1.5mu\overline{\mkern-1.5muR\mkern-1.5mu}\mkern 1.5mu}$ is decomposable, a contradiction. ∎
From Lemma 7.7 it follows that
$\mathcal{A}=\mathcal{A}_{W_{1}A}\wr_{(W_{1}A)/A}\mathcal{A}_{G/A}~\text{or}~\mathcal{A}=\mathcal{A}_{W_{1}H}\wr_{(W_{1}H)/A}\mathcal{A}_{G/A}.$
In the first case, $\mathcal{A}_{A}$ is $\otimes$-complemented in $\mathcal{A}_{W_{1}}$ by Eq. (16). So the section $(W_{1}A)/A$ satisfies Condition $(2)$ from Theorem 7.3 and we are done.
In the second case, $W_{1}H$ is an $\mathcal{A}$-subgroup. Let $a$ be a nontrivial element of $A$, $b\in H\setminus A$, and $T$ a basic set of $\mathcal{A}$ containing $b$. Clearly, $W_{1}H=W_{1}A\times\langle b\rangle$, $W_{1}A$ is an $\mathcal{A}_{W_{1}H}$-subgroup, and $W_{1}$ is the largest $\mathcal{A}_{W_{1}H}$-subgroup of odd order. Therefore the $S$-ring $\mathcal{A}_{W_{1}H}$, the basic set $T$, and the subgroup $W_{1}$ satisfy [14, Hypothesis $5.1$]. The structure of a basic set containing an involution of an $S$-ring satisfying [14, Hypothesis $5.1$] is described in [14, Lemma $5.5$]. Applying this lemma to $T$ and using the inclusion $\{a\}=A^{\#}\in\mathcal{S}(\mathcal{A})$ one can deduce that
$T=W_{0}Ab~\text{or}~T=W_{0}b$
for some $\mathcal{A}_{W_{1}}$-subgroup $W_{0}$. To complete the proof of the lemma, let us consider the above two cases separately.
Let $T=W_{0}Ab$. As $H=\langle Ab\rangle$ is not an $\mathcal{A}$-subgroup, $T\neq Ab$ and hence $|W_{0}|>1$. So $W_{0}$ satisfies the condition of [14, Lemma $5.4$]. Applying this lemma to $\mathcal{A}_{W_{1}H}$, we conclude that $\mathcal{A}_{W_{1}H}$ is the $(W_{1}A)/(W_{0}A)$-wreath product. Therefore $\mathcal{A}$ is the $(W_{1}A)/A$-wreath product. Again, the section $(W_{1}A)/A$ satisfies Condition $(2)$ from Theorem 7.3 by Eq. (16) and we are done.
Now let $T=W_{0}b$, then $W_{2}=\langle T,W_{1}\rangle$ is an $\mathcal{A}_{W_{1}H}$-subgroup such that $|W_{2}\cap A|=1$ and $W_{1}H=W_{2}\times A$. Lemma 3.2(2) implies that $\mathcal{A}_{W_{1}H}=\mathcal{A}_{W_{2}}\otimes\mathcal{A}_{A}$. Therefore $(W_{1}H)/A$ satisfies Condition $(2)$ from Theorem 7.3 as desired. ∎
Theorem 7.3 immediately follows from Propositions 7.4 and 7.5.
Theorem 7.8.